# On the Regularization of Learnable Embeddings for Time Series Processing

### v1

@@ 4.1 Regularization methods
L1 and L2 regulariation
L2 regularization (Krogh & Hertz, 1991), also known as weight decay, consists in adding to the loss a penalty term proportional to the magnitude of the model’s weights. When applied to the embeddings, the penalty term is $\mathcal{L}_{l2}\left({\bm{E}}\right)=\lambda_{l2}\cdot\|\mathbf{E}\|_{2}^{2}$. Similarly, L1 regularization (Tibshirani, 1996), also known as lasso, consists in applying a penalty term of $\mathcal{L}_{l1}\left({\bm{E}}\right)=\lambda_{l1}\cdot\|\mathbf{E}\|_{1}$. The positive scalar values, $\lambda_{l2}$ and $\lambda_{l1}$, control the regularization strength.
@@
The clustering regularization introduced in (Cini et al., 2023b) is based on learning a set of cluster centroids and a cluster assignment matrix, alongside the embeddings. A regularization term is then added to the loss to minimize the distance between embeddings and the assigned centroid. It was originally designed to structure the embedding space and improve model transferability.
@@
Variational regularization (Cini et al., 2023b), consists in modeling embeddings as samples from a Gaussian posterior learned from data $\mathbf{e}^{i}\sim\mathcal{N}\left(\bm{\mu}^{i},diag\left(\bm{\sigma}^{i}\right)\right)$, where $\bm{\mu}^{i}$ and $\bm{\sigma}^{i}$ are the learnable local parameters. A penalty term based on the KL-divergence between the learned distribution and a standard Gaussian prior, is added to the loss. Similarly to the clustering regularization, it was introduced to enhance model transferability.
@@
We also consider a novel forgetting regularization, based on the forget-and-relearn paradigm (Zhou et al., 2021). This refers to strategies where the parameters of some neural network layers are occasionally reset during training. By doing this, it is possible to reduce memorization of training samples (Baldock et al., 2021), avoiding shortcuts and learning representations that generalize better (Geirhos et al., 2020; Zhou et al., 2021). In practice, we periodically reset the local embeddings ${\bm{E}}$ to a sample from a shared initialization distribution $\mathcal{P}_{e}$ (Narkhede et al., 2022), every $K$ training epochs. Concurrently, we similarly reset the Encoder’s  and Decoder’s weights (Eq. 7) that directly multiply the embeddings (see Appendix C.1). The value of $K$ can be easily selected empirically (see Appendix C.2). Additionally, resetting is halted after a certain amount of epochs to allow for convergence to a final configuration. This can be tuned manually or automatically, by monitoring the validation loss.
@@ 4.2 Preventing co-adaptation
As mentioned in Sec. 1, jointly learning local and global parameters can result in their co-adaptation. In particular, embeddings might become simple sequence identifiers and result in models that rely entirely on this identification mechanism. Besides the negative effects on sample efficiency, the resulting architecture would likely lose flexibility in terms of the transferability of the learned representations and processing blocks. In the empirical analysis, we will then focus on whether methods designed to break this co-adaptation can be effective regularizations. Intuitively, the global processing block would be less likely to rely on embeddings as identifiers if these are actively perturbed during training, e.g., by resampling or zeroing.
Among the regularization methods considered in the analysis, L1 and L2 regularizations simply penalize the embeddings’ magnitude. Similarly, clustering encourages embeddings to occupy specific regions within the embedding space and does not actively perturb their values. Conversely, by randomly zeroing out embeddings’ parameters, we might expect dropout to prevent co-adaptation more effectively, as similarly happens when applying it to the weights of subsequent layers (Srivastava et al., 2014). The variational regularization can have a similar impact, as embeddings are constantly resampled during training. Finally, forgetting can prevent co-adaptation in the early stages of training, by periodically resetting the local parameters. Empirical evidence of the effectiveness of such design principles is provided in the following.
@@ 5.1 Time series forecasting benchmarks
Table 1: Forecasting test error under optimal model size and learning rate (5 runs, ±$\pm 1\sigma$). Methods equal or better than the corresponding standard architectures (RNN, STGNN, STAtt) with embeddings (+Emb.) are in bold. The best-performing method within each dataset and model is in red.
@@
In our first experiment, we consider the problem of time series forecasting in a transductive setting, i.e., the set of time series to be forecast is the same set observed at training time.
@@
The forecasting horizon is dataset-dependent and reported in Tab. 4 of Appendix A. We use models without local parameters and with un-regularized local embeddings as reference baselines. Then, we evaluate performance with regularized local embeddings, adopting the different strategies detailed in Sec. 4.1. For model selection, a hyperparameter search on model size and learning rate has been carried out independently for each model variant and dataset. All regularization hyperparameters have been set as detailed in Appendix D. Tab. 1 shows the obtained results, while Tab. 2 provides a concise summary of the performance of each regularization across experimental settings. Regularizing the learning of local embeddings provides performance improvements over non-regularized models in most datasets (F1). Considering the negligible computational overhead and ease of implementation, these results support the adoption of such techniques as standard practice. While there is no clear winner among the different regularization methods, the variational regularization appears to be the most effective, on average; followed by forgetting and dropout. This supports adopting methods that prevent co-adaptation by design, i.e., that prevent the global components from relying on specific values of the local parameters (F2) (F3). Additional results on the impact of regularizing embeddings when Eq. 4 is implemented by a simple MLP are reported in Appendix E.
@@ 5.2 Sensitivity analysis and learning curves
Figure 3: Validation curves for different training scenarios (5 runs, ±$\pm 1\sigma$). Plot names follow the convention model (embedding size) [dataset].
To provide additional insight, we investigate how different regularizations affect the learning curve across different embedding sizes. In doing so, we fix the shared model hidden size to $h=64$ and learning rate to $\eta=0.00075$. Specifically, Fig. 3 shows the validation mean absolute error (MAE) across training epochs for some example scenarios; different colors correspond to different regularization strategies. For completeness, additional plots in complementary settings are reported in Appendix F. Comparing the different regularizations, we can see that dropout (red) and forgetting (blue) significantly affect the learning curves, even with large embeddings (Fig. 3, right). However, dropout, as one could expect, might be problematic when applied to embeddings of limited size (Fig. 3, left). The forgetting regularization is robust across the different configurations. The variational regularization appears less disruptive of the learning dynamics. The remaining regularizations (i.e., L1, L2, and clustering) have little impact on the learning curve. Overall, regularization strategies that perturb the local embeddings seem more likely to positively affect the learning curves, which might be a reason for the superior performance shown in Tab. 2.
@@ 5.3 Transfer learning
Table 3: Forecasting test error (MAE) in transfer learning (5 runs, ±$\pm 1\sigma$, STGNN). Within each target dataset, different rows pertain to different fine-tuning data lengths. Bold denotes methods equal or better than the baseline (+ Emb.), best method is in red.
@@
Our second main experiment consists of time series forecasting under a transfer learning setting, adapted from previous works (Cini et al., 2023b). This benchmark aims to verify the impact of regularizing the local embeddings on the transferability of the global (shared) processing blocks. We consider the $4$ PEMS benchmark datasets from Guo et al. (2021) (i.e., PEMS03, PEMS04, PEMS07, PEMS08) and a reference STGNN model with local embeddings. For each subset of three datasets, we first train the entire model, then we reset the embeddings to their initial values and fine-tune them exclusively on the held-out dataset. During fine-tuning, shared parameters are kept frozen. The different regularizations are applied only during the initial training procedure and disabled during fine-tuning. Tab. 3 reports the results for the un-regularized and regularized models as the amount of data for fine-tuning varies (zero-shot refers to no fine-tuning at all). It is evident how dropout, variational regularization, and forgetting are the most reliable and effective strategies (F2). Notably, dropout excels in the zero-shot setting, while forgetting is consistent across fine-tuning lengths and excels when more data are available (F3). This suggests dropout can be particularly useful when data from the transfer domain is scarce or absent.
@@ 5.4 Robustness to local parameter perturbation
Figure 4: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1\sigma$, STGNN, METR-LA). Left: Adding zero-mean Gaussian noise. Right: (Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal.
Finally, we investigate how a model’s forecasting accuracy is affected by the perturbation of the learned local embeddings. This provides insights into the robustness of the shared learned parameters and can serve as a proxy for the effectiveness of different regularizations in preventing co-adaptation. To avoid penalizing regularizations that are sensitive to the weights’ magnitude  (i.e., L2, L1), we consider perturbations that do not impact the scale of the learned representations. In particular, we experiment with four such strategies: adding zero-mean Gaussian noise to the embeddings (Noise $\sigma$), randomly shuffling embeddings across sequences (Rearranged), replacing each embedding with their sample mean (Mean), resampling each embedding from their sample normal distribution (Sampled) (see Appendix D.1 for a formal definition). Fig. 4 shows the results of the analysis after training an STGNN on the METR-LA dataset (complementary settings are illustrated in Appendix F). Regularization methods that actively perturb the embeddings while learning (i.e., dropout, forgetting, variational) consistently result in more robust models under all the considered perturbations (F2) (F3). Conversely, other regularizations have marginal, or even negative, impact. We observe some consistency between the results observed in Fig. 4 and performance in transfer learning, shown in Tab. 3.
@@ 6 Conclusions
This paper highlights the importance of regularizing the learning of local embeddings in modern deep-learning architectures for time series forecasting. Our empirical study, across diverse datasets and scenarios, provides clear evidence that this practice is beneficial for a variety of reference architectures, representative of the state of the art. Notably, even simple techniques, when applied to local embeddings, can yield consistent performance gains. Furthermore, we observed that methods that actively perturb the embeddings at training time, such as dropout, variational regularization, and forgetting, consistently rank among the top performers, both in transductive and transfer learning settings. Overall, our findings offer a strong empirical argument for the adoption of embedding regularization as standard practice when designing hybrid global-local architectures for time series. We believe this is an important building block towards developing more robust and transferable models for spatiotemporal data and designing foundation models for related time series processing.
@@
While it is clear that these regularizations provide consistent advantages, it is difficult to strongly identify the best-performing method in the different scenarios. Hence, future works might focus on finding strategies that combine the best qualities of the different methods. Furthermore, future studies could try to provide methods to quantitatively and analytically characterize the co-adaptation of global and local parameters, for which we can only identify indirect signs. In particular, they might design experiments to detect the degeneration of local components into mere identifiers. Moreover, future research could explore how the efficacy of different regularization methods varies w.r.t. the number of input time series and different downstream tasks.
@@ Appendix B Models
$\mathbf{m}^{j\rightarrow i}$ $=\mathbf{W}_{2}\xi\left(\mathbf{W}_{1}\left[{\bm{h}}^{i}\|{\bm{h}}^{j}\|a_{ji}\right]\right),\qquad\alpha^{j\rightarrow i}=\sigma\left(\mathbf{W}_{0}\mathbf{m}^{j\rightarrow i}\right),$ (8)
$\mathbf{\tilde{h}}^{i}$ $=\xi\left(\mathbf{W}_{3}{\bm{h}}^{i}+\sum_{j\in\mathcal{N}\left(i\right)}\{\alpha^{j\rightarrow i}\mathbf{m}^{j\rightarrow i}\}\right),$ (9)
@@
$\mathbf{\tilde{h}}^{i}=\xi\left(\mathbf{W}_{1}{\bm{h}}^{i}+\frac{1}{\|\mathcal{N}\left(i\right)\|}\sum_{j\in\mathcal{N}\left(i\right)}\{\mathbf{W}_{2}{\bm{h}}^{j}\}\right),$ (10)

### v2

@@ 4.1 Regularization methods
L1 and L2 regularization
L2 regularization (Krogh & Hertz, 1991), also known as weight decay, consists in adding to the loss a penalty term proportional to the magnitude of the model’s weights. When applied to the embeddings, the penalty term is $\mathcal{L}_{l2}\left({\bm{E}}\right)=\lambda_{l2}\cdot\|{\bm{E}}\|_{2}^{2}$. Similarly, L1 regularization (Tibshirani, 1996), also known as lasso, consists in applying a penalty term of $\mathcal{L}_{l1}\left({\bm{E}}\right)=\lambda_{l1}\cdot\|{\bm{E}}\|_{1}$. The positive scalar values, $\lambda_{l2}$ and $\lambda_{l1}$, control the regularization strength.
@@
The clustering regularization introduced in (Cini et al., 2023b) is based on learning a set of cluster centroids and a cluster assignment matrix, alongside the embeddings. A regularization term, with weight $\lambda_{clst}$, is then added to the loss to minimize the distance between embeddings and the assigned centroid. It was originally designed to structure the embedding space and improve model transferability.
@@
Variational regularization (Cini et al., 2023b), consists in modeling embeddings as samples from a Gaussian posterior learned from data $\mathbf{e}^{i}\sim\mathcal{N}\left(\bm{\mu}^{i}_{var},diag\left(\bm{\sigma}^{i}_{var}\right)\right)$, where $\bm{\mu}^{i}_{var}$ and $\bm{\sigma}^{i}_{var}$ are the learnable local parameters. A penalty term based on the KL-divergence between the learned distribution and a standard Gaussian prior, with weight $\lambda_{var}$, is added to the loss. Similarly to the clustering regularization, it was introduced to enhance model transferability.
@@
We also consider a novel forgetting regularization, based on the forget-and-relearn paradigm (Zhou et al., 2021). This refers to strategies where the parameters of some neural network layers are occasionally reset during training. By doing this, it is possible to reduce memorization of training samples (Baldock et al., 2021), avoiding shortcuts and learning representations that generalize better (Geirhos et al., 2020; Zhou et al., 2021). We adopt this paradigm in the context of hybrid global-local time series forecasting architectures and propose the following procedure. Every $K$ training epochs, we periodically reset local embeddings ${\bm{E}}$ to a sample from a shared initialization distribution $\mathcal{P}_{e}$. Concurrently, we similarly reset the Encoder’s  and Decoder’s weights (Eq. 7) that directly multiply the embeddings (see Appendix C.1). The value of $K$ can be easily selected empirically (see Appendix C.2). Additionally, resetting is halted after a certain amount of epochs to allow for convergence to a final configuration; this can be easily tuned manually, by considering the model’s convergence speed or triggered automatically, by monitoring the validation loss. Note that resetting local parameters to regularize the training of a related time series forecasting model has never been explored in the literature.
@@ 4.2 Preventing co-adaptation
As mentioned in Sec. 4, learning local and global parameters jointly on a downstream task can result in overfitting. We use the term co-adaptation to indicate this phenomenon, by referring to global and local parameters co-adapting during training. Specifically, local embeddings might simply act as sequence identifiers and result in models that rely entirely on this identification mechanism. Besides the negative effects on sample efficiency, the resulting architecture would likely lose flexibility in terms of the transferability of the learned representations and processing blocks. Furthermore, this behavior is in contrast with the intended one for local embeddings, i.e., encoding meaningful characteristics of the target time series. One way in which regularization methods can aim to prevent such co-adaptation is to actively perturb the local parameters during training so that subsequent layers cannot rely on specific feature values (Srivastava et al., 2014). In our case, the global processing block would be less likely to rely on specific values of the embeddings when processing the target time series. In the following, we will use the term perturbation in a broad sense, referring to strategies that modify the values of local parameters at training time, e.g., resampling, zeroing, or adding noise.
Among the considered regularization methods, dropout perturbs the embeddings by randomly zeroing out parameters, as similarly done for the inputs/outputs of subsequent layers to avoid the co-adaptation of the associated weights (Srivastava et al., 2014). The variational regularization, given a common prior, learns each embedding’s sampling distribution end-to-end while resampling their actual values before providing them as inputs to the downstream at each forward pass. Finally, forgetting regularization introduces perturbation in the early stages of training by periodically resetting the local parameters. In particular, by considering this strategy for the first time in this context, we aim to test whether actively perturbing local parameters offers an effective principle for the design of new regularizations for hybrid time series forecasting models.
Regarding the other regularization methods considered in the analysis, L1 and L2 regularizations simply penalize the embeddings’ magnitude, while clustering similarly penalizes the distance of the embeddings from the learned centroids, forcing them to occupy specific regions within the embedding space. While these regularizations can provide structure to the embedding space, they do not actively perturb embeddings’ values. The empirical results presented in Sec. 5 altogether suggest that strategies based on embedding perturbations are more effective at preventing the kind of overfitting we previously discussed, i.e., co-adaptation.
@@ 5.1 Time series forecasting benchmarks
Table 1: Forecasting test error under optimal model size and learning rate (5 runs, ±$\pm 1std$). Methods equal to or better than the corresponding standard architectures (RNN, STGNN, STAtt) with embeddings (+ Emb.) are in bold. The best-performing method within each dataset and model is in red. ‘+ Regularization’ denotes the addition of that specific regularization only, on top of ‘+ Emb.’ .
@@
In our first experiment, we consider the problem of time series forecasting in a transductive setting, i.e., the set of time series to be forecast is the same set observed at training time. The forecasting horizon is dataset-dependent and reported in Tab. 4 of Appendix A. We use models without local parameters and with un-regularized local embeddings as reference baselines. Then, we evaluate performance with regularized local embeddings, adopting the different strategies detailed in Sec. 4.1. For model selection, a hyperparameter search on the hidden size (i.e., the number of units in each model’s hidden layer) and learning rate has been carried out independently for each model variant and dataset. All regularization hyperparameters have been set as detailed in Appendix D.
@@
Tab. 1 shows the obtained results, while Tab. 2 provides a concise summary of the performance of each regularization across experimental settings. Regardless of the strategy, regularizing the learning of local embeddings provides consistent performance improvements over non-regularized models in most datasets (F1). Considering the negligible computational overhead and ease of implementation, these results support the adoption of such techniques as standard practice. While there is no clear winner among the different regularization methods, the variational regularization appears to be the most effective, on average; followed by forgetting and dropout. This suggests that methods perturbing the embedding values can be more effective in preventing co-adaptation (F2) (F3). We can speculate that the effectiveness of the variational regularization stems from it jointly perturbing embeddings’ values while also providing structure to the embedding space. Additional results on the impact of regularizing embeddings when Eq. 4 is implemented by a simple MLP are reported in Appendix E, while results regarding the combination of multiple regularization techniques can be found in Appendix F.
@@ 5.2 Sensitivity analysis and learning curves
Figure 3: Validation curves for different training scenarios (5 runs, ±$\pm 1std$). Plot names follow the convention model (embedding size) [dataset].
To provide additional insight, we investigate how different regularizations affect the learning curve across different embedding sizes. In doing so, we fix the shared model hidden size to $d_{h}=64$ and learning rate to $lr=0.00075$. Specifically, Fig. 3 shows the validation mean absolute error (MAE) across training epochs for some example scenarios; different colors correspond to different regularization strategies. For completeness, additional plots in complementary settings are reported in Appendix G. Comparing the different regularizations, we can see that dropout (red) and forgetting (blue) significantly affect the learning curves, even with large embeddings (Fig. 3, right). However, dropout, as one could expect, might be problematic when applied to embeddings of limited size (Fig. 3, left). The forgetting regularization is robust across the different configurations. The variational regularization appears less disruptive of the learning dynamics. The remaining regularizations (i.e., L1, L2, and clustering) have little impact on the learning curve. Overall, regularization strategies that perturb the local embeddings seem more likely to positively affect the learning curves, which might be a reason for the superior performance shown in Tab. 2.
@@ 5.3 Transfer learning
Table 3: Forecasting test error (MAE) in transfer learning (5 runs, ±$\pm 1std$, STGNN). Within each target dataset, different rows pertain to different fine-tuning data lengths. Bold denotes methods equal to or better than the baseline (+ Emb.), best method is in red. ‘+ Regularization’ denotes the addition of that specific regularization only, on top of ‘+ Emb.’ .
@@
Our second main experiment consists of time series forecasting under a transfer learning setting, adapted from previous works (Cini et al., 2023b). This benchmark aims to verify the impact of regularizing the local embeddings on the transferability of the global (shared) processing blocks. We consider the $4$ PEMS benchmark datasets from Guo et al. (2021) (i.e., PEMS03, PEMS04, PEMS07, PEMS08) and a reference STGNN model with local embeddings. For each subset of three datasets, we first train the entire model, then we reset the embeddings to their initial values and fine-tune them exclusively on the held-out dataset. During fine-tuning, shared parameters are kept frozen. The different regularizations are applied only during the initial training procedure and disabled during fine-tuning. Tab. 3 reports the results for the un-regularized and regularized models as the amount of data for fine-tuning varies (zero-shot refers to no fine-tuning at all). It is evident how dropout, variational regularization, and forgetting are the most reliable and effective strategies (F2). Notably, dropout excels in the zero-shot setting, while forgetting is consistent across fine-tuning lengths and excels when more data are available (F3). This suggests dropout can be particularly useful when data from the transfer domain is scarce or absent. Conversely, regularizations based on structuring the embedding space, i.e., L1, L2, and the clustering regularization, show mixed results compared to the unconstrained model.
@@ 5.4 Robustness to local parameter perturbation
Figure 4: Test performance degradation on embeddings perturbation (5 runs, ±$\pm 1std$, STGNN, METR-LA). Left: Adding zero-mean Gaussian noise. Right: (Left) random shuffling, (Middle) replaced with their mean, and (Right) replaced by a draw from their sample normal.
Finally, we investigate how a model’s forecasting accuracy is affected by the perturbation of the learned local embeddings. This provides insights into the robustness of the shared learned parameters and can serve as a proxy for the effectiveness of different regularizations in preventing co-adaptation. To avoid penalizing regularizations that are sensitive to the weights’ magnitude  (i.e., L2, L1), we consider perturbations that do not impact the scale of the learned representations. In particular, we experiment with four such strategies: adding zero-mean Gaussian noise to the embeddings (Noise $\sigma$), randomly shuffling embeddings across sequences (Rearranged), replacing each embedding with their sample mean (Mean), resampling each embedding from their sample normal distribution (Sampled) (see Appendix D.1 for a formal definition). Fig. 4 shows the results of the analysis after training an STGNN on the METR-LA dataset (complementary settings are illustrated in Appendix G). As one might expect, regularization methods that actively perturb the embeddings while learning (i.e., dropout, forgetting, variational) consistently result in more robust models under all the considered perturbations (F2) (F3). Conversely, other regularizations have marginal, or even negative, impact. We observe some consistency between the results observed in Fig. 4 and performance in transfer learning, shown in Tab. 3.
5.5 Discussion
Considering the observed empirical results, we can summarize a set of practical recommendations that can provide practitioners with a starting point for model search. In transductive settings, variational regularization performs reliably across benchmarks and it is a safe choice. Moreover, we suggest favoring techniques that perturb the embedding values during training, e.g., variational regularization, forgetting or dropout. In the context of transfer learning, our results suggest the adoption of dropout, when fine-tuning the model on a low amount of data, while forgetting or variational regularization can give better results as more data become available. The recommendation of prioritizing techniques that perturb the embeddings at training time remains valid in the transfer learning settings.
@@ 6 Conclusions
This paper highlights the importance of regularizing the learning of local embeddings in modern deep-learning architectures for time series forecasting. Our empirical study, across diverse datasets and scenarios, provides clear evidence that this practice is beneficial for a variety of reference architectures, representative of the state of the art. Notably, even simple techniques, when applied to local embeddings, can yield consistent performance gains. Furthermore, we observed that methods that actively perturb the embeddings at training time, such as dropout, variational regularization, and forgetting, consistently rank among the top performers, both in transductive and transfer learning settings. The use of this property as a guideline in designing new regularization methods is further supported by the effectiveness of the unprecedentedly considered forgetting regularization. Overall, our findings offer a strong empirical argument for the adoption of embedding regularization as standard practice when designing hybrid global-local architectures for time series. We believe this is an important building block towards developing more robust and transferable models for spatiotemporal data and designing foundation models for related time series processing.
@@
While it is clear that these regularizations provide consistent advantages, it is difficult to strongly identify the best-performing method in the different scenarios. Hence, future works might focus on finding strategies that combine the best qualities of the different methods. Furthermore, future studies could try to provide methods to quantitatively and analytically characterize the co-adaptation of global and local parameters, for which we can only identify indirect signs. In particular, they might design experiments to detect the degeneration of local components into identifiers. Moreover, future research could explore how the efficacy of different regularization methods varies with the number of input time series and different downstream tasks.
Acknowledgments
Partly supported by International Partnership Program of the Chinese Academy of Sciences under Grant 104GJHZ2022013GC
@@ Appendix B Models
$\mathbf{m}^{j\rightarrow i}$ $=\mathbf{W}_{2}\xi\left(\mathbf{W}_{1}\left[{\bm{h}}^{i},{\bm{h}}^{j},a_{ji}\right]\right),\qquad\alpha^{j\rightarrow i}=\eta\left(\mathbf{W}_{0}\mathbf{m}^{j\rightarrow i}\right),$ (8)
$\mathbf{\tilde{h}}^{i}$ $=\xi\left(\mathbf{W}_{3}{\bm{h}}^{i}+\sum_{j\in\mathcal{N}\left(i\right)}\{\alpha^{j\rightarrow i}\odot\mathbf{m}^{j\rightarrow i}\}\right),$ (9)
@@
$\mathbf{\tilde{h}}^{i}=\xi\left(\mathbf{W}_{4}{\bm{h}}^{i}+\frac{1}{\|\mathcal{N}\left(i\right)\|}\sum_{j\in\mathcal{N}\left(i\right)}\{\mathbf{W}_{5}{\bm{h}}^{j}\}\right),$ (10)